OpenStudy (haley.fieldss):

Solve the equation. round to the nearest tenth if necessary. 2. 2x+12=18- x?

OpenStudy (gunsodo):

2x+12=18-x 2x+x=18-12 3x=6 x=2 :)

KjSaif (kjsaif):

Simplify both sides of the equation.

KjSaif (kjsaif):

2x+12=−x+18

KjSaif (kjsaif):

Add x to both sides.

KjSaif (kjsaif):

2x + 12 + x = −x + 18 + x

KjSaif (kjsaif):

3x+12=18

KjSaif (kjsaif):

Subtract 12 from both sides.

KjSaif (kjsaif):

3x+12−12=18−12

KjSaif (kjsaif):

3x=6

KjSaif (kjsaif):

Then divide both sides by 3.

OpenStudy (haley.fieldss):

So 2 is the correct answer?

KjSaif (kjsaif):

3x/3 = 6/3

KjSaif (kjsaif):

Yes! x=2

KjSaif (kjsaif):

Tag me if you need any help! :)